OpenStudy (anonymous):

The radius of the cone is 1.75 inches, and its height is 3.5 inches. If the diameter of the bubble gum ball is 0.5 inches, what is the closest approximation of the volume of the cone that can be filled with flavored ice?

OpenStudy (anonymous):

A snow cone is a tasty treat with flavored ice and a spherical bubble gum ball at the bottom.

OpenStudy (anonymous):

11.15 in3 11.22 in3 0.07 in3 33.67 in3

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

spherical bublegum

OpenStudy (anonymous):

volume of a sphere!!!!!!!!!!

OpenStudy (anonymous):

the volume of the bubble gum is .0654 approximately

OpenStudy (anonymous):

\[V_{cone} = \frac{1}{3}πr^{2}h\] \[V_{sphere} = \frac{4}{3}πr^{3}\] \[Answer = \frac{V_{cone}}{V_{sphere}}\]

OpenStudy (anonymous):

k

OpenStudy (anonymous):

now the volume of the cone

OpenStudy (anonymous):

the volume of the cone is 11.225

OpenStudy (anonymous):

I'm thinking about substraction right now

OpenStudy (anonymous):

11.15 looks promising

OpenStudy (anonymous):

Well, I'm confused. Why is the bubblegum even there?

OpenStudy (anonymous):

ehh... I don't know

OpenStudy (anonymous):

I'm assuming that the bubblegum's volume is .007

OpenStudy (anonymous):

11.22-.007=11.15

OpenStudy (anonymous):

And plus, Manjin Buu is taking over the world!!!!!!!!!!!!1

OpenStudy (anonymous):

If you have ever seen Dragonballz

OpenStudy (anonymous):

But don't you agree that 11.15 is the answer?

OpenStudy (anonymous):

I guess. I'm assuming that the cone has the bubblegum in it and you're wondering how much room is left for flavored ice?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

11.22 is the volume for the cone

OpenStudy (anonymous):

I'm assuming .007 is the volume for the sphere

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

11.22-.007=11.15

OpenStudy (anonymous):

agreed?

OpenStudy (anonymous):

or you think I am lacking somewhere?

OpenStudy (anonymous):

cal?

OpenStudy (anonymous):

I agree

OpenStudy (anonymous):

ehh